# Learning Theory for Kernel Bilevel Optimization

### v1

@@ 6 Conclusion and Perspectives
In this work, we established the first statistical generalization bounds for (KBO). This paper is a first step in providing generalization results for bilevel gradient-based methods in a non-parametric setting. We showcased the applicability of these bounds on the simplest algorithms and expect that our results can be readily extended to state-of-the-art methods such as those discussed in (Arbel & Mairal, 2022b; Ghadimi & Wang, 2018; Chen et al., 2021; Dagréou et al., 2022). The results concerning $U$-processes, however, are derived under the restrictive assumption of i.i.d. data. Relaxing such assumption to handle Markovian data is a promising direction for future work, as it would allow considering reinforcement learning applications where bilevel optimization methods have shown promising results (Nikishin et al., 2022).
Acknowledgements
This work was supported by the ANR project BONSAI (grant ANR-23-CE23-0012-01). EP thank AI Interdisciplinary Institute ANITI funding, through the French “Investments for the Future – PIA3” program under the grant agreement ANR-19-PI3A0004, Air Force Office of Scientific Research, Air Force Material Command, USAF, under grant numbers FA8655-22-1-7012. EP thank TSE-P, acknowledge support from ANR Chess, grant ANR-17-EURE-0010, ANR Regulia, support from IUF. EP and SV acknowledge support from ANR MAD. SV thanks PEPR PDE-AI (ANR-23-PEIA-0004) and the chair 3IA Côte d’Azur BOGL.
@@ Acknowledgements
Roadmap. We begin by presenting and establishing regularity properties of the objective functions in Appendix A. In Appendix B, we introduce the gradient estimators. Appendix C is dedicated to proving the boundedness and Lipschitz continuity of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$, along with local boundedness and Lipschitz properties of $\ell_{in}$, $\ell_{out}$ and their derivatives. The generalization results are provided in Appendix D. In Appendix E, we establish maximal inequalities for bounded and Lipschitz families of functions. Differentiability properties of the objectives are studied in Appendix F. Finally, Appendix G contains auxiliary technical lemmas used throughout the proofs.
[…]
Appendix A Regularity and Differentiability Results
A.1 Regularity of the objectives
@@
The following propositions establish differentiability of considered objectives. We defer their proof to Appendix F.
Proposition A.1 (Differentiability of $L_{in}$ and $L_{out}$).
@@
Under Assumptions (B), (C) and (A), for any $(\omega,h)\in{\mathbb{R}^{d}}\times\mathcal{H}$, the functions $L_{in}$, $L_{out}$ admit finite values at $(\omega,h)$, are jointly differentiable in $(\omega,h)$, with gradient given by:
@@
Similarly, the empirical estimates $\widehat{L}_{in}$ and $\widehat{L}_{out}$ admit finite values, and are differentiable with gradients admitting similar expressions as above with $\mathbb{P}$ and $\mathbb{Q}$ replaced by their empirical estimates ^$\hat{\mathbb{P}}_{n}$ and ^$\hat{\mathbb{Q}}_{m}$.
Proposition A.2 (Differentiability of $\partial_{h}L_{in}$).
Under Assumptions (B), (C) and (A), for any $(\omega,h)\in{\mathbb{R}^{d}}\times\mathcal{H}$, the functions $(\omega,h)\mapsto\partial_{h}L_{in}(\omega,h)$ is differentiable with partial derivatives given by:
@@
Moreover, for any $\omega\in\mathbb{R}^{d}$ and $h\in\mathcal{H}$, the operators $\partial_{\omega,h}^{2}L_{in}(\omega,h)$ and $\partial_{h}^{2}L_{in}(\omega,h)-\lambda\operatorname{Id}_{\mathcal{H}}$ are Hilbert-Schmidt, i.e., bounded operators with finite Hilbert-Schmidt norm. The same conclusions hold for the empirical estimate $(\omega,h)\mapsto\partial_{h}\widehat{L}_{in}(\omega,h)$ with partial derivatives admitting similar expressions as above with $\mathbb{P}$ replaced by its empirical estimate ^$\hat{\mathbb{P}}_{n}$.
Proposition A.3 (Strong convexity of the inner objective in its second variable and invertibility of the Hessians).
Under Assumptions (D), (B), (C) and (A), $h\mapsto L_{in}(\omega,h)$ and $h\mapsto\widehat{L}_{in}(\omega,h)$ are $\lambda$-strongly convex for any $\omega\in{\mathbb{R}^{d}}$. Moreover, for any $\omega\in\mathbb{R}^{d}$ and $h\in\mathcal{H}$, the Hessian operators $\partial_{h}^{2}L_{in}(\omega,h)$ and $\partial_{h}^{2}\widehat{L}_{in}(\omega,h)$ are invertible with their operator norm bounded by $\frac{1}{\lambda}$.
@@
By Assumption (D), we know that $v\mapsto\ell_{in}(\omega,v,y)$ is convex for any $\omega\in\mathbb{R}^{d}$ and $y\in\mathcal{Y}$. Moreover, by Proposition A.1, $(x,y)\mapsto\ell_{in}{\left(\omega,h(x),y\right)}$ is integrable for any $\omega\in\mathbb{R}^{d}$ and $h\in\mathcal{H}$. Consequently, by integration, we directly deduce that $h\mapsto\mathbb{E}_{\mathbb{P}}{\left[\ell_{in}{\left(\omega,h(x),y\right)}\right]}$ is convex for any $\omega\in{\mathbb{R}^{d}}$. Finally, $h\mapsto L_{in}(\omega,h)\coloneqq\mathbb{E}_{\mathbb{P}}{\left[\ell_{in}{\left(\omega,h(x),y\right)}\right]}+\frac{\lambda}{2}{\left\lVert h\right\rVert}_{\mathcal{H}}^{2}$ must be $\lambda$-strongly convex, for any $\omega\in\mathbb{R}^{d}$, as a sum of a convex function and a $\lambda$-strongly convex function. Similarly, we deduce that $h\mapsto\widehat{L}_{in}(\omega,h)$ is $\lambda$-strongly convex, for any $\omega\in\mathbb{R}^{d}$. Invertibility follows from the expression of the Hessian operator in Proposition A.2
A.2 Differentiability of the value function
Proposition A.4 (Total functional gradient $\nabla\mathcal{F}$).
Assume Assumptions (D), (A) and (C) hold. For any $\omega\in{\mathbb{R}^{d}}$, the total functional gradient $\nabla\mathcal{F}(\omega)$ satisfies:
@@ A.2 Differentiability of the value function
By applying Propositions A.1 and A.3, we know that $h\mapsto L_{in}(\omega,h)$ has finite values, is $\lambda$-strongly convex and Fréchet differentiable. Moreover, by Proposition A.2, $\partial_{h}L_{in}$ is Fréchet differentiable on ${\mathbb{R}^{d}}\times\mathcal{H}$, and, a fortiori, Hadamard differentiable. Therefore, by the functional implicit differentiation theorem (Ioffe & Tihomirov, 1979), we deduce that the map $\omega\mapsto h_{\omega}^{\star}$ is uniquely defined and is Fréchet differentiable with Jacobian $\partial_{\omega}h^{\star}_{\omega}$ solving the following linear system for any $\omega\in{\mathbb{R}^{d}}$:
@@
Using that $\partial_{h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is invertible by Proposition A.3, we can express $\partial_{\omega}h^{\star}_{\omega}$ as:
@@
Furthermore, $L_{out}$ is jointly Fréchet differentiable by application of Proposition A.1, so that $\omega\mapsto\mathcal{F}(\omega)$ is also differentiable by composition of the functions $(\omega,h)\mapsto L_{out}(\omega,h)$ and $\omega\mapsto(\omega,h^{\star}_{\omega})$. For a given $\omega\in{\mathbb{R}^{d}}$, the gradient of $\mathcal{F}$ is then given by the chain rule:
@@
Substituting the expression of $\partial_{\omega}h^{\star}_{\omega}$ into Equation 9 yields:
@@
To conclude, it suffices to notice that the function $a_{\omega}^{\star}$ appearing in Equation 7 must be equal to $-H_{\omega}^{-1}d_{\omega}$. Indeed, $a_{\omega}^{\star}$ is defined as the minimizer of the quadratic objective $L_{adj}(\omega,a)$ in Equation 8 which is strongly convex since the Hessian operator is lower-bounded by $\lambda\operatorname{Id}_{\mathcal{H}}$. Consequently, the minimizer $a_{\omega}^{\star}$ exists and is uniquely characterized by the optimality condition:
@@
Appendix B Gradient Estimators
Proposition B.1 (Expression of ^$\nabla\widehat{\mathcal{F}}(\omega)$ by implicit differentiation).
Under Assumptions (B), (D), (A) and (C), for any $\omega\in{\mathbb{R}^{d}}$, the gradient ^$\nabla\widehat{\mathcal{F}}(\omega)$ of the discretized functional bilevel optimization problem ($\widehat{\text{KBO}}$) is given by:
^$\nabla\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}-\frac{1}{m}\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}\mathbf{u}\in\mathbb{R}^{d},$
where $\operatorname{\mathbf{K}}$ and $\operatorname{\mathbf{\overline{K}}}$ are the Gram matrices in $\mathbb{R}^{n\times n}$ and $\mathbb{R}^{m\times n}$ with entries given by $\operatorname{\mathbf{K}}_{ij}\coloneqq K(x_{i},x_{j})$ and $\operatorname{\mathbf{K}}_{ij}\coloneqq K(\tilde{x}_{j},x_{j})$, and $\operatorname{\mathbf{M}}$, $\mathbf{u}$, $\operatorname{\mathbf{D_{v}^{out}}}$, $\operatorname{\mathbf{D_{v,v}^{in}}}$, $\operatorname{\mathbf{D_{\omega}^{out}}}$ and $\operatorname{\mathbf{D_{\omega,v}^{in}}}$ are defined as:
$\operatorname{\mathbf{M}}$ $\coloneqq\operatorname{\mathbf{K}}\operatorname{\mathbf{D_{v,v}^{in}}}+n\lambda\mathbbm{1}_{n\times n}{\in\mathbb{R}^{n\times n}},\qquad$ $\mathbf{u}$ $\coloneqq\operatorname{\mathbf{\overline{K}}}^{\top}\operatorname{\mathbf{D_{v}^{out}}}{\in\mathbb{R}^{n}},$
$\operatorname{\mathbf{D_{v}^{out}}}$ $\coloneqq{\left(\partial_{v}\ell_{out}{\left(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j}\right)}\right)}_{1\leq j\leq m}\in\mathbb{R}^{m},\quad$ $\operatorname{\mathbf{D_{\omega}^{out}}}$ $\coloneqq{\left(\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})\right)}_{1\leq j\leq m}\in\mathbb{R}^{d\times m},$
$\operatorname{\mathbf{D_{v,v}^{in}}}$ $\coloneqq\operatorname{\mathbf{diag}}{\left({\left(\partial_{v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\right)}_{1\leq i\leq n}\right)}\in\mathbb{R}^{n\times n},\qquad$ $\operatorname{\mathbf{D_{\omega,v}^{in}}}$ $\coloneqq{\left(\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\right)}_{1\leq i\leq n}\in\mathbb{R}^{d\times n}.$

### v3

@@ 6 Conclusion and perspectives
Summary. In this work, we established the first generalization bounds for (KBO). These results are crucial for understanding the generalization properties of algorithms for solving (KBO). They offer rigorous guarantees on the algorithm’s performance on unseen data—a fundamental criterion for any algorithmic design—and help control overfitting. Given that our bounds are of order $\mathcal{O}(\nicefrac{{1}}{{\sqrt{m}}}+\nicefrac{{1}}{{\sqrt{n}}})$, this highlights the equal importance of both outer- and inner-level sample sizes to the overall generalization error. Our findings can impact current practices, particularly in hyperparameter optimization, where the validation dataset is typically much smaller than the training set.
Limitations and future work. This paper takes a first step toward providing generalization results for bilevel gradient-based methods in a nonparametric setting. While our theoretical analysis focused on a full-batch bilevel setting with exact gradients, extending this framework to stochastic variants, such as those in [Arbel:2021a, ghadimi2018approximation, chen2021closing, Dagreou2022SABA], remains an open challenge. A promising direction would be to consider approximate kernel representations, such as random Fourier features or neural tangent kernels, which enable scalable learning using kernel methods while preserving useful theoretical properties. Furthermore, the constants in our bounds are likely conservative; we did not investigate their tightness or potential for improvement. A deeper analysis of their optimality could provide valuable insights and constitutes an avenue worth exploring. Additionally, providing generalization guarantees for the un-regularized problem, possibly in the form of minimax optimal rates for (KBO), is a worthwhile future direction. This requires controlling the constants as $\lambda\to 0$, provided additional source assumptions are made, as discussed in [smale2007learning, sriperumbudur2017density]. Finally, broadening our framework to cover non-smooth losses, such as the hinge loss in SVMs, is an interesting direction for future work.
Acknowledgments and Disclosure of Funding
This work was supported by the ANR project BONSAI (grant ANR-23-CE23-0012-01). EP acknowledges funding from the AI Interdisciplinary Institute ANITI through the French Investments for the Future – PIA3 program under grant agreement ANR-19-PI3A-0004; the Air Force Office of Scientific Research, Air Force Materiel Command, USAF under grant number FA8655-22-1-7012; TSE-P; the ANR project Chess (grant ANR-17-EURE-0010); the ANR project Regulia; and the Institut Universitaire de France (IUF). EP and SV acknowledge support from the ANR project MAD. SV also thanks the PEPR PDE-AI program (ANR-23-PEIA-0004) and the 3IA Côte d’Azur Chair BOGL.
@@
Roadmap. In Appendix˜A, we review existing non-kernel generalization bounds in the bilevel optimization literature and explain why minimax optimal rates cannot be obtained in our setting. We begin the theoretical appendices by presenting and establishing regularity properties of the objective functions in Appendix˜B. In Appendix˜C, we introduce the gradient estimators. Appendix˜D is dedicated to proving the boundedness and Lipschitz continuity of $h^{\star}_{\omega}$ and $\hat{h}_{\omega}$, along with local boundedness and Lipschitz properties of $\ell_{in}$, $\ell_{out}$, and their derivatives. The generalization results are provided in Appendix˜E. In Appendix˜F, we establish maximal inequalities for bounded and Lipschitz families of functions. Differentiability properties of the objectives are studied in Appendix˜G. Appendix˜H contains auxiliary technical lemmas used throughout the proofs. Finally, further details on the experiments and additional numerical results are provided in Appendix˜I.
Notations. $\|\cdot\|$ denotes the Euclidean norm in $\mathbb{R}^{d}$, $\|\cdot\|_{\mathcal{H}}$ denotes the norm in the RKHS $\mathcal{H}$, $\|\cdot\|_{\operatorname{op}}$ denotes the operator norm, and $\|\cdot\|_{\operatorname{HS}}$ denotes the Hilbert-Schmidt norm. $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ denotes the inner product on $\mathcal{H}$, and $\langle\cdot,\cdot\rangle_{\operatorname{HS}}$ denotes the Hilbert-Schmidt inner product. $K(x,\cdot)$ denotes the feature map, for any $x\in\mathcal{X}$. For any two normed spaces $E$ and $F$, $\mathcal{L}(E,F)$ denotes the space of continuous linear operators from $E$ to $F$. For any two probability distributions $\mathcal{P}$ and $\mathcal{Q}$, $\mathcal{P}\otimes\mathcal{Q}$ denotes the product measure of $\mathcal{P}$ and $\mathcal{Q}$. Given two Hilbert spaces $\left(H_{1},\langle\cdot,\cdot\rangle_{H_{1}}\right)$ and $\left(H_{2},\langle\cdot,\cdot\rangle_{H_{2}}\right)$, the tensor product of $u\in H_{1}$ and $v\in H_{2}$, denoted by $u\otimes v$, is an operator from $H_{2}$ to $H_{1}$ defined, for any $e\in H_{2}$, as $\left(u\otimes v\right)e=u\left\langle v,e\right\rangle_{H_{2}}$. For any $v_{1},\ldots,v_{n}\in\mathbb{R}$, $\operatorname{\mathbf{diag}}(v_{1},\ldots,v_{n})\in\mathbb{R}^{n\times n}$ denotes a diagonal matrix of size $n\times n$, where the diagonal entries are $v_{1},\ldots,v_{n}$ and all the off-diagonal entries are 0. $\mathbbm{1}_{m}$ denotes a vector of size $m$ where all entries are $1$. $\mathbbm{1}_{n\times n}$ denotes the identity matrix of size $n$. For any vector space $V$ over $\mathbb{R}$, $\operatorname{Id}_{V}$ denotes the identity operator on $V$. Given a compact set $\mathcal{K}$, $\operatorname{diam}(\mathcal{K})$ denotes its diameter. $v^{\top}$ denotes the transpose of either a vector or a matrix, depending on the context.
Appendix A Further Discussion
Existing generalization bounds for the non-kernel case. bao2021stability laid foundational work towards understanding generalization in bilevel optimization by analyzing uniform stability in full-batch bilevel optimization. Their generalization criterion compares the population outer loss evaluated at the output of a randomized algorithm to the empirical outer loss evaluated using the same algorithm. Given a number $\kappa$ between 0 and 1, they obtain a decay at a rate of $\mathcal{O}(\nicefrac{{T^{\kappa}}}{{m}})$ for unrolled optimization, which decreases as $\nicefrac{{1}}{{m}}$ in outer sample size, but increases with the number of outer iterations $T$ made. This criterion differs from ours, which instead compares the population outer objective at the theoretically optimal inner solution $h^{\star}_{\omega}$ to the empirical loss evaluated at the empirical solution $\hat{h}_{\omega}$. Complementing this upper bound, wang2024lower established lower bounds on the uniform stability of gradient-based bilevel algorithms, demonstrating a rate of $\Omega(\nicefrac{{1}}{{m}})$. Building on [bao2021stability], zhang2024fine extended the analysis to stochastic bilevel optimization, establishing on-average stability bounds and deriving a generalization rate of $\mathcal{O}(\nicefrac{{1}}{{\sqrt{m}}})$ due to the presence of stochastic gradients. In a related context, oymak2021generalization studied generalization in neural architecture search using a bilevel formulation, showing that approximate inner solutions and Lipschitz continuity of the outer loss yield a generalization bound of $\mathcal{O}(\nicefrac{{1}}{{\sqrt{m}}}+\nicefrac{{1}}{{\sqrt{n}}})$. arora2020provable investigated representation learning for imitation learning via bilevel optimization, offering generalization bounds of order $\mathcal{O}(\nicefrac{{1}}{{\sqrt{m}}})$ that depend both on the size of the dataset and the stability of learned representations.
Difficulty of obtaining minimax rates in our setting. Although spectral filtering yields minimax rates [meunier2024nonparametric] in the kernel instrumental variable regression setting [singh2019kernel], it fundamentally relies on a linear operator representation of the inner minimizer, typically characterized through the spectral decomposition of a compact, self-adjoint covariance operator (see, e.g., [CaponnettoDeVito2007, bauer2007regularization]). This formulation allows one to apply functional calculus on the spectrum, with filter functions (such as Tikhonov regularization and truncated SVD) controlling the contribution of small eigenvalues [engl1996regularization, LoGerfoRosascoOdoneDeVitoVerri2008]. Under suitable source conditions, this enables the derivation of minimax optimal convergence rates for kernel ridge regression and other problems involving quadratic losses. In our framework, however, the inner objective is generally not quadratic in $\omega$, and the mapping $\omega\mapsto h^{\star}_{\omega}$ is nonlinear. Moreover, we consider a fixed $\lambda$, in contrast to the vanishing-regularization regimes where spectral filtering is most effective for minimax analysis. As a result, the source condition assumption and the spectral filtering tools that underpin minimax guarantees in the quadratic case do not apply directly in our setting.
Appendix B Regularity and Differentiability Results
B.1 Regularity of the objectives
@@
The following propositions establish differentiability of considered objectives. We defer their proof to Appendix˜G.
Proposition B.1 (Differentiability of $L_{in}$ and $L_{out}$).
@@
Under Assumptions˜(A), (C), (D) and (B), for any $(\omega,h)\in{\mathbb{R}^{d}}\times\mathcal{H}$, the functions $L_{in}$ and $L_{out}$ admit finite values at $(\omega,h)$, are jointly differentiable in $(\omega,h)$, with gradients given by:
@@
Similarly, the empirical estimates $\widehat{L}_{in}$ and $\widehat{L}_{out}$ admit finite values, and are differentiable with gradients admitting similar expressions as above with $\mathbb{P}$ and $\mathbb{Q}$ replaced by their empirical estimates $\widehat{\mathbb{P}}_{n}$ and $\widehat{\mathbb{Q}}_{m}$.
Proposition B.2 (Differentiability of $\partial_{h}L_{in}$).
Under Assumptions˜(A), (C), (D) and (B), for any $(\omega,h)\in{\mathbb{R}^{d}}\times\mathcal{H}$, the function $(\omega,h)\mapsto\partial_{h}L_{in}(\omega,h)$ is differentiable with partial derivatives given by:
@@
Moreover, for any $\omega\in\mathbb{R}^{d}$ and $h\in\mathcal{H}$, the operators $\partial_{\omega,h}^{2}L_{in}(\omega,h)$ and $\partial_{h}^{2}L_{in}(\omega,h)-\lambda\operatorname{Id}_{\mathcal{H}}$ are Hilbert-Schmidt, i.e., bounded operators with finite Hilbert-Schmidt norm. The same conclusions hold for the empirical estimate $(\omega,h)\mapsto\partial_{h}\widehat{L}_{in}(\omega,h)$ with partial derivatives admitting similar expressions as above with $\mathbb{P}$ replaced by its empirical estimate $\widehat{\mathbb{P}}_{n}$.
Proposition B.3 (Strong convexity of the inner objective in its second variable and invertibility of the Hessians).
Under Assumptions˜(A), (E), (C), (D) and (B), $h\mapsto L_{in}(\omega,h)$ and $h\mapsto\widehat{L}_{in}(\omega,h)$ are $\lambda$-strongly convex for any $\omega\in{\mathbb{R}^{d}}$. Moreover, for any $\omega\in\mathbb{R}^{d}$ and $h\in\mathcal{H}$, the Hessian operators $\partial_{h}^{2}L_{in}(\omega,h)$ and $\partial_{h}^{2}\widehat{L}_{in}(\omega,h)$ are invertible with their operator norm bounded by $\frac{1}{\lambda}$.
@@
By Assumption˜(E), we know that $v\mapsto\ell_{in}(\omega,v,y)$ is convex for any $\omega\in\mathbb{R}^{d}$ and $y\in\mathcal{Y}$. Moreover, by Proposition˜B.1, $(x,y)\mapsto\ell_{in}\left(\omega,h(x),y\right)$ is integrable for any $\omega\in\mathbb{R}^{d}$ and $h\in\mathcal{H}$. Consequently, by integration, we directly deduce that $h\mapsto\mathbb{E}_{\mathbb{P}}\left[\ell_{in}\left(\omega,h(x),y\right)\right]$ is convex for any $\omega\in{\mathbb{R}^{d}}$. Finally, $h\mapsto L_{in}(\omega,h)\coloneqq\mathbb{E}_{\mathbb{P}}\left[\ell_{in}\left(\omega,h(x),y\right)\right]+\frac{\lambda}{2}\left\lVert h\right\rVert_{\mathcal{H}}^{2}$ must be $\lambda$-strongly convex, for any $\omega\in\mathbb{R}^{d}$, as a sum of a convex function and a $\lambda$-strongly convex function. Similarly, we deduce that $h\mapsto\widehat{L}_{in}(\omega,h)$ is $\lambda$-strongly convex, for any $\omega\in\mathbb{R}^{d}$. Invertibility follows from the expression of the Hessian operator in Proposition˜B.2
B.2 Differentiability of the value function
Proposition B.4 (Total functional gradient $\nabla\mathcal{F}$).
Assume Assumptions˜(A), (E), (B) and (D) hold. For any $\omega\in{\mathbb{R}^{d}}$, the total functional gradient $\nabla\mathcal{F}(\omega)$ satisfies:
@@ B.2 Differentiability of the value function
By applying Propositions˜B.1 and B.3, we know that $h\mapsto L_{in}(\omega,h)$ has finite values, is $\lambda$-strongly convex and Fréchet differentiable. Moreover, by Proposition˜B.2, $\partial_{h}L_{in}$ is Fréchet differentiable on $\mathbb{R}^{d}$$\times\mathcal{H}$, and, a fortiori, Hadamard differentiable. Therefore, by the functional implicit differentiation theorem [ioffe1979theory, petrulionyte2024functional, Theorem 2.1], we deduce that the map $\omega\mapsto h_{\omega}^{\star}$ is uniquely defined and is Fréchet differentiable with Jacobian $\partial_{\omega}h^{\star}_{\omega}$ solving the following linear system for any $\omega\in{\mathbb{R}^{d}}$:
@@
Using that $\partial_{h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is invertible by Proposition˜B.3, we can express $\partial_{\omega}h^{\star}_{\omega}$ as:
@@
Furthermore, $L_{out}$ is jointly Fréchet differentiable by application of Proposition˜B.1, so that $\omega\mapsto\mathcal{F}(\omega)$ is also differentiable by composition of the functions $(\omega,h)\mapsto L_{out}(\omega,h)$ and $\omega\mapsto(\omega,h^{\star}_{\omega})$. For a given $\omega\in{\mathbb{R}^{d}}$, the gradient of $\mathcal{F}$ is then given by the chain rule:
@@
Substituting the expression of $\partial_{\omega}h^{\star}_{\omega}$ into Equation˜8 yields:
@@
To conclude, it suffices to notice that the function $a_{\omega}^{\star}$ appearing in Equation˜6 must be equal to $H_{\omega}^{-1}d_{\omega}$. Indeed, $a_{\omega}^{\star}$ is defined as the minimizer of the quadratic objective $L_{adj}(\omega,a)$ in Equation˜7 which is strongly convex since the Hessian operator is lower-bounded by $\lambda\operatorname{Id}_{\mathcal{H}}$. Consequently, the minimizer $a_{\omega}^{\star}$ exists and is uniquely characterized by the optimality condition:
@@
Appendix C Gradient Estimators
Proposition C.1 (Expression of $\nabla\widehat{\mathcal{F}}(\omega)$ by implicit differentiation).
Under Assumptions˜(C), (E), (B) and (D), for any $\omega\in{\mathbb{R}^{d}}$, the gradient $\nabla\widehat{\mathcal{F}}(\omega)$ of the discretized kernel bilevel optimization problem ($\widehat{\text{KBO}}$) is given by:
$\nabla\widehat{\mathcal{F}}(\omega)=\frac{1}{m}\operatorname{\mathbf{D_{\omega}^{out}}}\mathbbm{1}_{m}-\frac{1}{m}\operatorname{\mathbf{D_{\omega,v}^{in}}}\operatorname{\mathbf{M}}^{-1}\mathbf{u}\in\mathbb{R}^{d},$
where $\operatorname{\mathbf{K}}$ and $\operatorname{\mathbf{\overline{K}}}$ are the Gram matrices in $\mathbb{R}^{n\times n}$ and $\mathbb{R}^{m\times n}$ with entries given by $\operatorname{\mathbf{K}}_{ij}\coloneqq K(x_{i},x_{j})$ and $\operatorname{\mathbf{\overline{K}}}_{ij}\coloneqq K(\tilde{x}_{i},x_{j})$, and $\operatorname{\mathbf{M}}\in\mathbb{R}^{n\times n}$, $\mathbf{u}\in\mathbb{R}^{n}$, $\operatorname{\mathbf{D_{v}^{out}}}\in\mathbb{R}^{m}$, $\operatorname{\mathbf{D_{\omega}^{out}}}\in\mathbb{R}^{d\times m}$, $\operatorname{\mathbf{D_{v,v}^{in}}}\in\mathbb{R}^{n\times n}$, and $\operatorname{\mathbf{D_{\omega,v}^{in}}}\in\mathbb{R}^{d\times n}$ are defined as:
$\operatorname{\mathbf{M}}$ $\coloneqq\operatorname{\mathbf{K}}\operatorname{\mathbf{D_{v,v}^{in}}}+n\lambda\mathbbm{1}_{n\times n},\qquad$ $\mathbf{u}$ $\coloneqq\operatorname{\mathbf{\overline{K}}}^{\top}\operatorname{\mathbf{D_{v}^{out}}},$
$\operatorname{\mathbf{D_{v}^{out}}}$ $\coloneqq\left(\partial_{v}\ell_{out}\left(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j}\right)\right)_{1\leq j\leq m},\quad$ $\operatorname{\mathbf{D_{\omega}^{out}}}$ $\coloneqq\left(\partial_{\omega}\ell_{out}(\omega,\hat{h}_{\omega}(\tilde{x}_{j}),\tilde{y}_{j})\right)_{1\leq j\leq m},$
$\operatorname{\mathbf{D_{v,v}^{in}}}$ $\coloneqq\operatorname{\mathbf{diag}}\left(\left(\partial_{v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\right)_{1\leq i\leq n}\right),\qquad$ $\operatorname{\mathbf{D_{\omega,v}^{in}}}$ $\coloneqq\left(\partial_{\omega,v}^{2}\ell_{in}(\omega,\hat{h}_{\omega}(x_{i}),y_{i})\right)_{1\leq i\leq n}.$